OpenStudy (anonymous):

|x-6| is less than or equal to -12. How do i solve this?

OpenStudy (anonymous):

there are no solutions to this

OpenStudy (anonymous):

how can the absolute value less than or equal to a negative no?

OpenStudy (anonymous):

\[|x-6|\le -12\]

OpenStudy (anonymous):

Thats what it says, hahaha. Im as confused as you are!

OpenStudy (anonymous):

Recall that: \[|a| \le b \iff -b \le a \le b\] Therefore \[|x-6| \le -12 \implies 12 \le x-6 \le -12\] Which has no solutions, because nothing is both less than -12 and greater than 12.

OpenStudy (amistre64):

i assume that these problems are designed to test that you comprehend what is going on and not just simply trying to follow a mechanical step to get an answer. There are many illogical ways to put together a sentence in math; just as there is in speech

OpenStudy (anonymous):

It was just an empty set, haha. My bad!